Hi what are the next two numbers in the sequence;
15.2, 11.64, 8.16

Answer by greenestamps(13334)



The question is invalid; there is no way to know the answer.

Use any two numbers you want.

If there is no information about what kind of sequence this is supposed to be, then any next two numbers will make a valid sequence.

One clue to how the sequence MIGHT HAVE BEEN formed is the ratio or difference between successive terms. In this sequence...

15.2-11.64 = 3.56
11.64-8.16 = 3.48

The difference decreased by 0.08; IF that is the pattern for the sequence, then

the 4th term is 8.16-3.40 = 4.76
the 5th term is 4.76-3.32 = 1.44

But, in the end, WHATEVER pattern you find is only a guess; the actual sequence might have been formed by a completely different (or even random) process.